How long does it take for a ball dropped off of a building of height 69.2 m to hit the ground if the ball starts it's fall from rest?


1
Expert's answer
2021-04-09T11:25:43-0400

Answer

By equation of motion

"S=ut+\frac{1}{2}gt^2"


Initially u=0m/s

So

"69.2=\frac{1}{2}(9.81) t^2\\t=3.75s"